# Improving Relational Database Interactions with Large Language Models: Column Descriptions and Their Impact on Text-to-SQL Performance

### v1

@@ B.1.2 Annotating the Quality of the Improved Column Descriptions
B.1.3 Annotating the Difficulty of Predicting the Column Descriptions for the BIRD Dataset
@@
B.1.4 Annotating the Quality of Generated Column Descriptions
@@
Somewhat correct The column description is somewhat correct but has room for improvement due to missing information. For example, "The Time column records the specific time at which a transaction occurred, formatted in a 24-hour $H$:$M$ pattern," which lacks enough information to make a valid prediction beyond the primary purpose.
@@
B.2 Annotation Scores
Descriptions Agreement Percentage Cohen’s Kappa
Original Descriptions 0.79 0.68
Improved Descriptions 0.82 0.61
Difficulty Annotation Agreement Percentage Cohen’s Kappa
Difficulty Rating 0.53 0.31
B.3 Examples of Columns, Gold Descriptions and Difficulties
@@
Easy superhero gender gender The name of the gender of the superhero. Self explanatory.
Easy student_club event event_date The date the event took place or is scheduled to take place. Self explanatory.
Medium student_club zip_code short_state Two-letter state abbreviation. While short_state suggests state abbreviations, it requires looking at the data to confirm it represents two-letter state codes, thus needing some context.
Medium formula_1 results laps The number of laps completed by the driver in a race. Requires understanding of the racing domain, intuitive for basic knowledge of sports.
Hard card_games foreign_data uuid Unique identifier referencing the related card. Concept of a UUID may not be clear, and connection to "related card" needs understanding of data relationships.
Hard formula_1 results rank The overall rank of the driver’s fastest lap time in the race. Without context, rank could refer to different types of rankings (e.g., race position, qualifying rank), making it ambiguous without more information.
Very Hard financial district A4 The number of inhabitants in the district. Column name provides no intuitive clue. Data only contains numbers giving no clue as to the nature of the data.
Very Hard financial district A11 Numeric code for the district. The column name A11 in the Districts table is uninformative and, with only numeric data, it’s impossible to discern its purpose without documentation.
B.4 Prompt Templates
@@
7""
8{database_schema}
9""
@@
11Hereisexampledatafromthetable{table}:""
12
13{example_data}
14
15Hereisupto10possibleuniquevaluesforthecolumn{column}fromthetable
16{table}:
17
18{unique_data}
@@
20""
21
22Thecolumnnamefor{column}is{column_name}.Thisisthenameofthecolumn,
23itcancontainimportantinformationaboutthecolumn,andshouldbeusedto
24writethedescription.
25Thepreviouscolumndescriptionis{column_description}.Thisistheold
26descriptionofthecolumn,thisissometimeslackingandshouldbereadand
27rewritten.
28
29###Task
30Generateaprecisedescriptionforthe{column}columninthe{table}table.
31Yourdescriptionshouldinclude:
32-Primarypurposeofthecolumn.Ifthedetailsintheschemadonotsuffice
33toascertainwhatthedatais,return:"Not enough information to make a
34valid prediction."
35Optionally,yourdescriptioncouldalsoinclude:
36-Additionalusefulinformation(ifapparentfromtheschema),formatted
37asanewsentence,butnevermorethanone.Ifnousefulinformationis
38availableorifthedetailsintheschemadonotsufficetoascertain
39usefuldetails,returnnothing.
40
41###Requirements
42-Focussolelyonconfirmeddetailsfromtheprovidedschema.
43-Keepthedescriptionconciseandfactual.
44-Excludeanyspeculativeoradditionalcommentary.
45-DONOTreturnthephrase"in the {table} table"inyourdescription.
46Thisveryimportant.
47
48**Examples:**
49-Foracolumnnamed"no. of municipalities with inhabitants < 499,"
50thedescriptionshouldbe:"This is the number of municipalities with
51fewer than 499 inhabitants."
52-Foracolumnitisbettertodescribedasshortaspossible:So,
53"Details about the ratio of urban inhabitants."ispreferredover"This
54column provides information on details about ratio of urban inhabitants,".
55-Ifthenameis"Frequency",thedescriptionshouldbe:" The frequency
56of transactions on the account"sincethiscomesfromtheaccounttable.
57-Ifthenameis"amount of money"thedescriptionsshouldbe:"
58The amount of money in the order"sincethiscomesfromtheordertable.
59
60###Pleaseskipthe"data_format"andfocussolelyonupdatingthe
61"column_description".
62
63DONOTreturnanythingelseexceptthegeneratedcolumndescription.
64Thisisveryimportant.Theanswershouldbeonlythegeneratedtext
65aimedatdescribingthecolumn.
66"""
Figure 5: The prompt used to improve upon BIRD-Bench already existing column descriptions.

### v3

@@ B.7.2 Annotating the Quality of the Improved Column Descriptions
B.7.3 Annotating the Difficulty of Predicting the Column Descriptions for the BIRD Dataset
@@
B.7.4 Annotating the Quality of Generated Column Descriptions
@@
Somewhat correct The column description is somewhat correct but has room for improvement due to missing information. For example, "The Time column records the specific time at which a transaction occurred, formatted in a 24-hour HH:MM pattern," which lacks enough information to make a valid prediction beyond the primary purpose.
@@
B.8 Examples of Columns, Gold Descriptions and Difficulties
@@
Easy superhero gender gender The name of the gender of the superhero. Self-explanatory.
Easy student_club event event_date The date the event took place or is scheduled to take place. Self-explanatory.
Medium student_club zip_code short_state Two-letter state abbreviation. Requires context: while short_state suggests state abbreviations, it needs data confirmation to represent two-letter state codes.
Medium formula_1 results laps The number of laps completed by the driver in a race. Requires basic knowledge of racing and sports.
Hard card_games foreign_data uuid Unique identifier referencing the related card. Understanding the concept of a UUID and its relation to the card requires some domain knowledge.
Hard formula_1 results rank The overall rank of the driver’s fastest lap time in the race. Without context, rank could refer to various rankings, making it ambiguous without more info.
Very Hard financial district A4 The number of inhabitants in the district. Column name provides no intuitive clue. Data contains only numbers, offering no indication of its purpose.
Very Hard financial district A11 Numeric code for the district. The column name A11 in the Districts table is uninformative; without documentation, it is impossible to infer its meaning.
B.9 Prompt Templates
@@
7"{database_schema}"
8
9Hereisexampledatafromthetable{table}:"{example_data}"
@@
11Hereisupto10possibleuniquevaluesforthecolumn{column}fromthe
12table{table}:{unique_data}
13
14Thecolumnnamefor{column}is{column_name}.Thisisthenameofthecolumn,
15itcancontainimportantinformationaboutthecolumn,andshouldbeusedto
16writethedescription.Thepreviouscolumndescriptionis{column_description}.
17Thisistheolddescriptionofthecolumn,thisissometimeslackingandshould
18bereadandrewritten.
@@
20###Task
21Generateaprecisedescriptionforthe{column}columninthe{table}table.
22Yourdescriptionshouldinclude:
23-Primarypurposeofthecolumn.Ifthedetailsintheschemadonotsuffice
24␣␣toascertainwhatthedatais,return:"Not enough information to make a valid
25 prediction."
26Optionally,yourdescriptioncouldalsoinclude:
27-Additionalusefulinformation(ifapparentfromtheschema),formatted
28␣␣asanewsentence,butnevermorethanone.Ifnousefulinformationis
29␣␣available,returnnothing.
30
31###Requirements
32-Focussolelyonconfirmeddetailsfromtheprovidedschema.
33-Keepthedescriptionconciseandfactual.
34-Excludeanyspeculativeoradditionalcommentary.
35-DONOTreturnthephrase"in the {table} table"inyourdescription.
36
37**Examples:**
38-"no. of municipalities with inhabitants < 499":
39␣␣"This is the number of municipalities with fewer than 499 inhabitants."
40-Short:"Details about the ratio of urban inhabitants."
41-"Frequency":"The frequency of transactions on the account."
42-"amount of money":"The amount of money in the order."
43
44###Skipthe"data_format"andfocussolelyonupdatingthe
45"column_description".
46
47DONOTreturnanythingelseexceptthegeneratedcolumndescription.
48"""